find the slope of the line given 3x- 4y = 8
Answer by stanbon(75887)

find the slope of the line given
3x- 4y = 8
---
Solve for "y":
4y = 3x-8
y = (3/4)x -2
---
slope = 3/4
================
Cheers,
Stan H.
==============